OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (10, 35) and (150, 55), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.

OpenStudy (anonymous):

flvs

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation y = x - 3 using the slope-intercept method?

OpenStudy (anonymous):

Using complete sentences, explain how to graph x + 3y = 9 using the x- and y-intercepts method.

OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation x - 2y = 4 using the x- and y- intercepts method ? Plot the y-intercept (0, -2). Plot the intercept (4, 0). Count up 1 and right 2 from the y-intercept. Draw a line through the two points.

OpenStudy (anonymous):

can someoneee help

OpenStudy (anonymous):

what module and exam is it

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

its a lesson

OpenStudy (anonymous):

r u gunna help or not

OpenStudy (anonymous):

yeah which test is it i aready did it

OpenStudy (anonymous):

it module four quiz

OpenStudy (anonymous):

4. what

OpenStudy (anonymous):

theres a lot of test in module four type it out like 4.01 4.02 4.03 etc

OpenStudy (anonymous):

04.04

OpenStudy (anonymous):

yeah or watver

OpenStudy (anonymous):

ya there

OpenStudy (anonymous):

fine byw

OpenStudy (anonymous):

The formula for finding slope is y2-y1/x2-x1. If you put in the numbers up above (10,35) & (150,55) in this formula you will get the slope. 150-10/55-35= 140/20= 7/1= 7